# Object-centric proto-symbolic behavioural reasoning from pixels

### v2

@@ 3.2 Reasoning and control
OBR’s generative model allows it to infer the current state of the world, as well as predict future states that would arise as the result of the agent’s actions. In a (model-based) RL framework, we could now define rewards and learn a value function on which to base a behavioral policy. Here, we take a slightly different approach, which is seminal in the proposed architecture, based on the influential computational neuroscience framework of Active Inference [39, 33]. It posits that agents, through their actions, aim to minimize a variational free energy with respect to a probability distribution over states that they prefer to find themselves in [40, 41, 42]. This preference distribution entails a biased (conditional) prior on environmental states, which acts as an attractor for planning future behavior. Once this preference (or soft goal) is learnt the rest of the architecture machinery will drive the system (agent-world) to generate object-centric actions that minimize the discrepancy of the current and the desired state.
@@
We assume that the preference distribution is given by $\tilde{p}(\{{{\bm{\mathbf{s}}}^{\dagger}}_{T>t}^{(k)}\}_{k\in 1:K}|\{{\bm{\mathbf{\lambda}}}^{(k)}_{t}\}_{k\in 1:K})=\prod_{k}\mathcal{N}(\tilde{{\bm{\mathbf{\mu}}}}^{(k)},{\tilde{{\bm{\mathbf{\sigma}}}}{{}^{(k)}}^{2}})$, where $T$ is some time horizon and $t$ denotes the present. Note that this distribution is conditioned on the agent’s internal beliefs, via the variational parameters. We implement this mapping through a set-structured MLP network $\phi$ (architectural details in Appendix 6.1), which preserves the order of objects from input to output, and is invariant to the number of objects:
@@
Given a preference distribution $\tilde{p}({{\bm{\mathbf{s}}}^{\dagger}})$ [^4] Note that in this notation, we do not condition the preference on the current state belief as above, as the planning procedure described here can be applied to any preference distribution., the linearized dynamics model enables the agent to plan actions efficiently in closed form (without the need of rollouts). Specifically, the action plan ${\bm{\mathbf{\pi}}}=\begin{bmatrix}{\bm{\mathbf{a}}}_{t+1},\ \ldots,\ {\bm{\mathbf{a}}}_{t+T}\end{bmatrix}^{\text{T}}$ is computed by minimizing the path integral of the variational free energy [41] over some future time horizon (similarly to model predictive control) [^5] Note that the notation in this section omits the object index $k$ for legibility, as actions can be planned independently between objects (a strength of our approach).:
@@
Importantly, this optimization can be performed in closed form, by computing the discrepancy between the preferred state and the sequence of states that will unfold if no action is taken, and projecting this discrepancy (${\bm{\mathbf{e}}}$) onto the pseudo-inverse of a matrix ${\bm{\mathbf{U}}}$ that maps actions within the planning horizon to their (cumulative) effects over that time window. ${\bm{\mathbf{L}}}=\text{diag}(\tilde{{\bm{\mathbf{\sigma}}}}{{}^{(k)}}^{-2})$ is a diagonal matrix containing the precisions (inverse variances) of the preference distributions, and $\lambda_{a}$ controls the strength with which actions are regularized (shrunk) towards to a zero-mean prior. Further details are provided in Appendix 7.1.
@@ 4.1 Environment
We developed, Active dSprites, which is an “activated” version of the various multi-dSprites datasets that have been used in previous work on object-based visual inference (e.g. [10, 12]). Not only does it include (continuous) dynamics, but these dynamics can can be acted on by an agent (through continuous control actions that accelerate the objects). Thus, active dSprites is an interactive environment, rather than a dataset. We implemented two different scenarios in the active dSprites environment. The first, objects are 2.5-D shapes (e.g., squares, ellipses and hearts)—they have no depth dimension of their own, but can occlude each other within the depth dimension of the image. The second objects are 3D shapes (e.g., boxes, half-torus, cones) that can move in a $10\times 10\times 10$ m cubic space. These objects are much more complex to learn than the 2D version, as they have lighting and shadows, producing color gradients that need to be encoded. To randomize the environment, when an active dSprites instance is intialized, object shapes, positions, sizes and colors are all sampled uniformly at random. Initial velocities are drawn from a Normal distribution. Shape colors are sampled at discrete intervals spanning the full range of RGB-colors, while a background color is drawn from a set of evenly spaced grayscale values between black and white. Shapes are presented in random depth order. In the 3D environment the lighting sources are fixed. Code for the active dSprites environment can be found at github.com/neuro-ai-robotics/OBR.
@@ 4.2 Problem complexity analysis through baselines comparison
We evaluated the problem solving complexity of a basic conditional rule ”If Heart then move Boxes to the Right, otherwise move Ellipses to the Left” comparing different oracle and baseline algorithms with our approach 3. We evaluated different algorithms with full observability (i.e., with access to object locations): Linear-Quadratic Regulator control [43, 44], (implemented in the Python 3 control systems library), Soft-Actor-Critic (SAC) [45] and Proximal Policy Optimization (PPO) [46] (both used from the stable baselines 3 library [47]), and our OBR with perfect access to the true environmental state). And with partial observability (i.e., the agent only has access to the RGB image (pixels): SAC, PPO and OBR. For RL algorithms we gave dense rewards using the object distance between the current state and the desired state. For the OBR we only allow actions after frame 5. Here objects are 2.5 dimensional and during training the shape, color and positions are randomized and velocities are sampled from a Normal distribution with mean 0.
@@
These results first show, in line with previous works [21], that manipulating objects in the environment cannot be solved through non object-centric approaches, and highlights an important weakness of current pixel-based RL baselines. Second, it shows that a basic reasoning rule that implies interaction with the environment is already sufficiently complex. Thus, learning multiple rules and composition in this active dSprite synthetic environment is relevant.
@@ 4.2.1 Abstract behavioural reasoning with conditional rules
Figure 4: Behavioural reasoning with 2D objects. (a) Two instances of the active 2D dSprites solved by the OBR agent which has learnt the “IfHeart” conditional rule: ((((Heart $\to$ Squares to top-right $\land$ Ellipses to middle-right)(¬$\land$($\neg$) ∧ ( ¬ Heart $\to$ Squares to top-left $\land$ Ellipses to middle-left $\land$ Hearts to down-left)))). Environment row shows the real scenario with the objects and Imagined goal row show the desired preference of the agent. The red arrows are the control actions applied to the objects in every frame. The agent is passive until the 5th frame and the preference network waits three frames to provide an output (this allows perception to be stable). The true goal is the idealized desired location of the objects. (b) Two instances of the active 2D dSprites solved by the OBR agent which has learnt the Heart XOR Square task. (c and d) Statistical evaluation of both tasks accuracy (MSE with respect to the idealized true goal) for different number of objects in the scene. The training was performed with 3 objects. Colored lines describe the the MSE between ground-truth goals and actual object positions and velocities. Error bars reflect +/- 1 SEM.
@@
Figure 5: Behavioural reasoning with 3D objects. (a) Two instances of the active 3D dSprites solved by the OBR agent which has learnt the “IfHalfTorus” conditional rule: ((((HalfTorus $\to$ Boxes to top-right-middle $\land$ Cones to middle-right-middle)(¬$\land$($\neg$) ∧ ( ¬ HalfTorus $\to$ Boxes to top-left-front $\land$ Cones to middle-left-front $\land$ HalfTorus to down-left-front)))). Note that in the second instance there are only Boxes and HalfTorus and no Cones but the OBR solves the conditional task without troubles. Environment row shows the real scenario with the 3D objects and Imagined goal row show the desired preference of the agent. The red arrows are the control actions applied to the objects in every frame. The agent is passive until the 5th frame and the preference network waits three frames to provide an output (this allows perception to be stable). The true goal is the idealized desired location of the objects. (b) Statistical evaluation of the task accuracy (MSE with respect to the idealized true goal) for different number of objects in the scene. Blue lines describe the mean and the error bars denote the standard deviation. (c) Instances of the OBR agent sequential behaviour with an IfHalfTorus rule with different number of objects in the scene. The goal represents the ideal location of the objects given the rule. The desired goals of the agent (preferent internal state) are not shown and are learnt through unsupervised learning using the preference network.
@@ 4.2.2 Adaptation to changes in the environment
Figure 6: Adaptation. (a) Two instances of the active dSprites solved by the OBR agent which has learnt the “IfHeart” conditional rule: ((((Heart $\to$ Squares to top-right $\land$ Ellipses to middle-right)(¬$\land$($\neg$) ∧ ( ¬ Heart $\to$ Squares to top-left $\land$ Ellipses to middle-left $\land$ Hearts to down-left)))). At 4th frame one of the objects (in this case the heart is substituted by a square, forcing not only to change the object-centric perception but the conditional behaviour, as the resulting reasoning depends on the appearance of the Heart. (b) Statistical performance evaluation of 512 instances of the object substitution experiment. The blue line is the median and error bars define the inter-quartile range to deal with the skewed data due to failures.
@@ 4.2.3 Logical composition emergence
Figure 7: Logical composition. The OBR agent preference reasoning module is trained with two distinct rules and then tested in execution if the composed rule emerges. Task accuracy over 512 instances shows that OBR is combining both rules in the test phase.

### v3

@@ 3.2 Reasoning and control
OBR’s generative model allows it to infer the current state of the world, as well as predict future states that would arise as the result of the agent’s actions. In a (model-based) RL framework, we could now define rewards and learn a value function on which to base a behavioral policy. Here, we take a slightly different approach, which is seminal in the proposed architecture, based on the influential computational neuroscience framework of Active Inference [parr2022active, lanillos2021active]. It posits that agents, through their actions, aim to minimize a variational free energy with respect to a probability distribution over states that they prefer to find themselves in [van2020deep, millidge2021whence, fountas2020deep]. This preference distribution entails a biased (conditional) prior on environmental states, which acts as an attractor for planning future behavior. Once this preference (or soft goal) is learnt the rest of the architecture machinery will drive the system (agent-world) to generate object-centric actions that minimize the discrepancy of the current and the desired state.
@@
[…]
We implement this mapping through a set-structured MLP network $\phi$ (architectural details in Appendix 6.1), which preserves the order of objects from input to output, and is invariant to the number of objects:
@@
Given a preference distribution $\tilde{p}({{\bm{\mathbf{s}}}^{\dagger}})$ [^5] Note that in this notation, we do not condition the preference on the current state belief as above, as the planning procedure described here can be applied to any preference distribution., the linearized dynamics model enables the agent to plan actions efficiently in closed form (without the need of rollouts). Specifically, the action plan ${\bm{\mathbf{\pi}}}=\begin{bmatrix}{\bm{\mathbf{a}}}_{t+1},\ \ldots,\ {\bm{\mathbf{a}}}_{t+T}\end{bmatrix}^{\text{T}}$ is computed by minimizing the path integral of the variational free energy [millidge2021whence] over some future time horizon (similarly to model predictive control) [^6] Note that the notation in this section omits the object index $k$ for legibility, as actions can be planned independently between objects (a strength of our approach).:
@@
[…]
@@ 4.1 Environment
We developed, Active dSprites, which is an “activated” version of the various multi-dSprites datasets that have been used in previous work on object-based visual inference (e.g. [Greff2019a_Iodine, locatello2020SlotAttention]). Not only does it include (continuous) dynamics, but these dynamics can be acted on by an agent (through continuous control actions that accelerate the objects). Thus, active dSprites is an interactive environment, rather than a dataset. We implemented two different scenarios in the active dSprites environment. The first, objects are 2.5-D shapes (e.g., squares, ellipses and hearts)—they have no depth dimension of their own, but can occlude each other within the depth dimension of the image. The second objects are 3D shapes (e.g., boxes, half-torus, cones) that can move in a $10\times 10\times 10$ m cubic space. These objects are much more complex to learn than the 2D version, as they have lighting and shadows, producing color gradients that need to be encoded. To randomize the environment, when an active dSprites instance is intialized, object shapes, positions, sizes and colors are all sampled uniformly at random. Initial velocities are drawn from a Normal distribution. Shape colors are sampled at discrete intervals spanning the full range of RGB-colors, while a background color is drawn from a set of evenly spaced grayscale values between black and white. Shapes are presented in random depth order. In the 3D environment the lighting sources are fixed. Code for the active dSprites environment can be found at github.com/neuro-ai-robotics/OBR.
@@ 4.2 Problem complexity analysis through baselines comparison
We evaluated the problem solving complexity of a basic conditional rule ”If Heart then move Boxes to the Right, otherwise move Ellipses to the Left” comparing different oracle and baseline algorithms with our approach 3. We evaluated different algorithms with full observability (i.e., with access to object locations): Linear-Quadratic Regulator control [kalman1960contributions, kirk2004optimal], (implemented in the Python 3 control systems library), Soft-Actor-Critic (SAC) [haarnoja2018soft] and Proximal Policy Optimization (PPO) [schulman2017proximal] (both used from the stable baselines 3 library [raffin2021stable]), and our OBR with perfect access to the true environmental state). And with partial observability (i.e., the agent only has access to the RGB image (pixels): SAC, PPO and OBR. For RL algorithms we gave dense rewards using the object distance between the current state and the desired state. For the OBR we only allow actions after frame 5. Here objects are 2.5 dimensional and during training the shape, color and positions are randomized and velocities are sampled from a Normal distribution with mean 0.
@@
These results first show, in line with previous works [veerapaneni2020entity], that manipulating objects in the environment cannot be solved through non object-centric approaches, and highlights an important weakness of current pixel-based RL baselines. Second, it shows that a basic reasoning rule that implies interaction with the environment is already sufficiently complex. Thus, learning multiple rules and composition in this active dSprite synthetic environment is relevant.
4.3 Behavioural reasoning
@@ 4.3.1 Abstract behavioural reasoning with conditional rules
[…]
@@
[…]
@@ 4.3.2 Adaptation to changes in the environment
[…]
@@ 4.3.3 Logical composition emergence
Figure 7: Logical composition. The OBR agent’s preference reasoning module is trained on two rules separately and evaluated at test time on scenes where both conditions apply. For Rule 1, the agent learns to move Squares to the top-left when a Heart is present, and to the middle-right otherwise. For Rule 2, it learns to move Ellipses to the middle-left when a Heart is present, and to the middle-right otherwise. At test time, the agent must combine both mappings. The composed behavior, in which Squares move to the top-left and Ellipses to the middle-left (else, to the middle-right), is not provided during training and therefore emerges during inference. Task accuracy over 512 instances, measured as the MSE with respect to the ideal target locations, confirms that the agent recovers the composed rule during execution.
4.4 Systematic Ablation Study and Comparison
To understand how different components contribute to object-centric behavioral reasoning, we decompose OBR into three functional blocks: perception (object representation learning), dynamics (temporal prediction), and action (control generation). We evaluate each block in isolation to quantify its role and to clarify how specific architectural choices influence overall performance.
4.4.1 Perceptual module
A central design choice in object-centric perception is the form of inference used to obtain object representations. Iterative methods refine these representations through multiple optimization steps per frame, enabling adaptive correction of early estimates and the integration of information across time. Approaches such as the proposed OBR, OP3 [veerapaneni2020entity], and IODINE [Greff2019a_Iodine] belong to this family of iterative amortized inference methods. Fully amortized alternatives compute object representations in a single feedforward pass and prioritize computational efficiency. The dominant trend in recent literature extends this amortized direction, particularly through Slot Attention [locatello2020SlotAttention] and its variants.
We compare our iterative perceptual backbone, which extends the IODINE architecture for video prediction and is referred to as IODINE-dyn, to the fully amortized SAVi-dyn [kipfconditional_SAVi] across three dimensions: reconstruction quality, temporal prediction, and perceptual adaptation. Models were evaluated on 1000 test sequences from the Active DSprites 2D environment using three standard metrics: LPIPS for perceptual reconstruction quality, mIoU for pixel-level segmentation accuracy, and FG-ARI for clustering consistency of foreground object assignments (Table 2). Across these global metrics, IODINE-dyn achieves consistently higher performance, supporting the effectiveness of iterative inference under dynamic conditions.
Beyond aggregate metrics, the models differ in how their inference mechanisms respond to challenging visual situations. In our dataset, several factors such as small object size, muted colors, weak contrast, and partial occlusions are common, making some objects difficult to extract from a single frame. Iterative inference enables IODINE-dyn to progressively refine uncertain hypotheses as more evidence becomes available, often resolving initially ambiguous shapes over successive frames. Fully amortized inference, as used in SAVi, tends to produce stable representations but is less able to revise early commitments, which can lead to incomplete or imprecise segmentations when cues remain weak throughout a sequence. This pattern is reflected in the per-frame trends in Fig. 8, where IODINE-dyn improves over rollout length while SAVi-dyn remains comparatively stationary.
Table 2: Quantitative results for slot dynamics prediction and unsupervised object discovery. Visual quality of predicted frames is measured by LPIPS (lower is better), while object discovery quality is measured by mIoU and FG-ARI (higher is better). Best results are in bold.
Method LPIPS $\downarrow$ mIoU $\uparrow$ FG-ARI $\uparrow$
SAVi (8f) 0.206 $\pm$ .014 0.309 $\pm$ .019 0.573 $\pm$ .041
SAVi (16f) 0.222 $\pm$ .014 0.576 $\pm$ .016 0.425 $\pm$ .034
IODINE-dyn (ours) 0.054 $\pm$ .012 0.632 $\pm$ .034 0.758 $\pm$ .033
Quantitative results do not fully capture differences in shape fidelity. As shown in Fig. 8, SAVi often produces representations that encode color and approximate location but do not preserve object morphology, which limits its ability to distinguish between objects with similar appearance. IODINE-dyn, in contrast, maintains sharper and more structured boundaries across frames. This distinction is functionally relevant: in conditional behavioral reasoning tasks, identifying the correct object shape is required to ground the appropriate rule, and reduced shape fidelity can therefore hinder downstream decision making.
Figure 8: Challenging perceptual cases. We present a representative subset of sequences from the unsupervised object discovery task that are naturally difficult for object-centric models due to factors such as small object size, color similarity, low contrast, and partial occlusions. Rows correspond to ground-truth frames, IODINE-dyn and SAVi reconstructions using 8-frame and 16-frame refinement, respectively. In these conditions, SAVi often prioritizes color over geometry, leading to incorrect slot assignment, accidental merging or splitting of objects, or temporally stable yet inaccurate representations. Occlusions further amplify these errors, causing the model to miss partially hidden shapes or to misplace low-contrast objects, often producing blurred or inconsistent estimates. IODINE-dyn (ours) performs iterative refinement at each timestep. Ambiguous shapes gradually emerge as evidence accumulates, and object identity is preserved more consistently across frames. Across these examples, the figure illustrates how the models behave under challenging visual conditions and highlights the qualitative differences in their treatment of object geometry and occlusion.
Because SAVi often fails to recover object morphology beyond coarse Gaussian-like blobs, a shape-based substitution analysis is not informative for this model class. To assess adaptation under conditions where both methods can maintain object identity, we instead performed a color-based intervention. At frame 2, one object was replaced by an identical shape with a different color, and we measured how accurately each model adapted to this change using a color fidelity score, defined as the normalized Euclidean distance in RGB space (0 to 1, lower is better). This setup parallels the object substitution experiment but isolates color as the causal factor.
Across many sequences, IODINE-dyn adapts rapidly to the intervention, with error decreasing immediately after the change and continuing to improve over time. SAVi adapts more slowly. The 8-frame variant shows moderate improvement, while the 16-frame variant exhibits more variability across slots and often fails to consistently propagate the new appearance. We include both variants to illustrate the trade-off between temporal averaging and representational stability within fully amortized inference. Normalized RGB-distance values provide a quantitative reference for this pattern: immediately after the intervention (frame 1), the errors are 0.067±0.081 for IODINE-dyn, 0.268±0.110 for SAVi-8f, and 0.443±0.135 for SAVi-16f; by frame 3 they decrease to 0.037±0.045, 0.202±0.071, and 0.299±0.188, respectively. Overall, these results show that iterative refinement provides a clear advantage in handling abrupt changes in appearance, complementing its benefits for recovering shape-based distinctions (Fig. 9).
Figure 9: Color intervention. To evaluate adaptation to sudden changes in appearance in the unsupervised object discovery setting, one object is replaced at frame 2 by an identical shape of a different color. The figure presents multiple sequences showing ground-truth frames, IODINE-dyn reconstructions, and the 8-frame and 16-frame variants of SAVi. IODINE-dyn (ours) adapts quickly and maintains stable representations, while SAVi adjusts more slowly and may introduce inconsistencies across slots when integrating the new color. Quantitative RGB-distance measurements follow the same pattern and support the qualitative differences observed here.
4.4.2 Dynamics learning
A central question in object-centric reasoning is how temporal dependencies are encoded. To disentangle perception from dynamics, we evaluate models using ground-truth object states as input. This upper-bound analysis isolates the contribution of the dynamics model itself and removes perceptual noise and control-feedback effects.
We systematically compared four classes of dynamics architectures that span the main paradigms in structured object-centric learning: linear, recurrent, sparse-gated, and attention-based models. Linear approaches assume constant-coefficient state transitions and provide an interpretable baseline for temporal evolution. Recurrent architectures (GRU, LSTM) capture nonlinear dependencies either per object or jointly across the full system. Sparse-gated variants, inspired by event-driven mechanisms such as GateL0RD [gumbsch2021sparsely], selectively activate updates to improve temporal credit assignment. Attention-based dynamics (transformer variants) [locatello2020SlotAttention] introduce relational reasoning across objects through learned cross-object interactions.
In the full OBR architecture, an action network infers object-centric control signals from pixel-space action fields, and a goal network generates desired states conditioned on context. The dynamics component itself uses a second-order linear state-space transition with Gaussian noise, a standard assumption in state estimation. The inverse mapping from actions to state updates is learned with a neural network.
To evaluate these mechanisms, we organized the analysis into two experimental groups. Group A (Fig. 10) examines dynamics quality through multi-step prediction, permutation equivariance, and robustness to action noise. Group B (Fig. 11) assesses reasoning and generalization through counterfactual, interventional, and opposite-rule tests. Together, these experiments characterize how each architecture captures temporal structure, relational dependencies, and causal disentanglement in the object-centric state space.
Across permutation equivariance, noise robustness, and temporal extrapolation, the OBR linear state-space model shows consistently strong behaviour, delivering competitive performance across all tests while maintaining near-zero equivariance error (Fig. 10). In contrast, recurrent baselines show the anticipated tendency to accumulate prediction error over longer horizons, and attention-based models demonstrate no consistent benefit despite higher complexity. Overall, OBR dynamics delivers reliable accuracy with substantially fewer parameters, making it an appealing choice when latency and stability matter.
Figure 10: Dynamics quality. All dynamics models were trained on 8,000 trajectories of 50 steps each (400k state–action transitions) generated under the environment’s Newtonian physics. Episodes begin from randomized initial conditions (positions uniformly sampled, velocities drawn from a zero-mean Gaussian, and fixed object identities), while actions correspond to uniformly sampled accelerations that drive deterministic kinematic updates. This training setup provides broad state–action coverage without policy bias and isolates the quality of the learned transition models. The figure reports mean squared error for three evaluations: permutation equivariance, noise robustness, and temporal extrapolation. The permutation test assesses whether predictions remain consistent when the order of objects in the state representation is permuted and then restored. The noise test measures prediction stability under small perturbations to the control inputs during short teacher-forced rollouts. The extrapolation test evaluates multi-step autoregressive prediction accuracy over longer horizons. Across these diagnostics, the OBR linear state-space dynamics achieve near-perfect permutation equivariance and competitive performance on noise robustness and extrapolation, while several recurrent models show greater sensitivity to object ordering and error accumulation at longer horizons.
Under online adaptation during 10-step rollouts (one gradient update per prediction), the OBR dynamics model attains among the lowest errors across three out-of-distribution tests that probe whether the models learned the underlying Newtonian physics rather than memorized training correlations. The opposite-rule test swaps object shape labels while leaving physical states unchanged; since shape does not influence motion, low error indicates that the model learned pure kinematics independent of task-specific labels. The counterfactual test exchanges two objects’ initial positions and velocities, evaluating whether the model can reason over alternative initial conditions. The interventional test freezes one object’s state (fixed position and zero velocity), probing whether the model respects the independence of object dynamics. Heavier recurrent and attention-based architectures often require careful tuning for online stability and show more variable behavior, whereas the linear state-space dynamics provide a reliable and lightweight option for adaptive closed-loop settings.
Figure 11: Reasoning and generalization with online adaptation. All dynamics models were trained under identical conditions on 8,000 trajectories of 50 steps generated from the environment’s Newtonian physics. During evaluation, models perform one gradient update per timestep to adapt their predictions online. The figure reports mean squared error for three out-of-distribution tests that probe whether models learned the underlying physical dynamics rather than memorized training correlations. The opposite-rule test swaps object shape labels while keeping physical states unchanged; the counterfactual test exchanges the initial states of two objects; and the interventional test freezes one object’s motion. Across these evaluations, the OBR linear dynamics model maintains competitive accuracy, while larger attention-based models often require careful tuning and exhibit more variable behaviour.
4.4.3 Action module
Finally, we compared OBR’s action module and closed-form control (inspired by active inference-based planning  [lanillos2021active]) against the current monolithic state-of-the-art for visual world models with control, V-JEPA 2-AC  [assran2025v]. Both algorithms were evaluated on the conditional rule $(\text{A}\to\text{B})\land(\neg\text{A}\to\text{C})$, referred to as the If-Heart task.
To avoid bias from implementation-specific training procedures in the native V-JEPA loss formulation, we assess three actor-head variants of V-JEPA 2-AC: the base model, a Dataset Aggregation (DAgger) variant, and a behavioral cloning (BC) variant. All V-JEPA variants share the same frozen encoder–predictor backbone and were trained under identical baseline conditions (complete architectural and training configurations are reported in Appendix 6.3).
Figure 12 reports the evolution of prediction error over rollout length. OBR oracle, trained with oracle demonstrations (i.e., oracle-provided accelerations, identical to the supervision used by V-JEPA), achieves near-perfect performance (MSE $\approx$ 0.00001). The behavioral cloning variant of V-JEPA (BC) reaches MSE $\approx$ 0.021 under identical training conditions. Standard OBR achieves MSE $\approx$ 0.013 using self-generated demonstrations from its action and preference modules. The performance difference reflects architectural distinctions. OBR employs object-based representations and modular components (action and preference modules) that generalize across different conditional rules with the same trained mechanism. In contrast, the V-JEPA BC variant exhibits convergence in this rule, but requires oracle-provided actions (accelerations) during training. More broadly, V-JEPA operates on monolithic pixel-level representations and requires oracle demonstrations for each new rule.
[…]